### v3

@@ 8.2 Results
Figure 5: Actual adoptions by the agency using the platform and other channels compared to Florida AFCARS benchmark model.
The agency has shared other benefits of implementing the platform. First, it enabled caseworkers to work more efficiently by avoiding the process of asking families to respond to announcements and saved time and emotional energy for all participants. Other programs, such as the diligent recruitment program described by Feldman et al. (2016), that showed improvement in children’s outcomes required advanced skills and high levels of labor from caseworkers, who are part of a workforce suffering from high burnout and turnover amidst increasing pressures on Florida’s child welfare system. With technology like the platform that we study, overburdened caseworkers could reallocate time spent reading dozens or hundreds of home studies from prospective families who expressed interest in a particular child. Second, the platform also helped to protect the privacy of the children in need of adoptive placements. Without the caseworker-driven search practice, their status as children in need of adoption would have been announced much more widely.
8.3 Validation from State Reports
To validate this analysis — especially to understand how the circuit that implemented the platform compares to statewide averages before and after implementation — we use the “Adoption Incentive” annual reports published by the Florida Department of Children and Families (2019, 2023). The analysis using AFCARS data implicitly assumes that the agency is representative of statewide patterns; if the agency already outperformed statewide averages before 2018, the analysis would overestimate the effect of the platform. The Adoption Incentive reports provide annual statistics for how the 20 circuits in Florida perform on various measures. Of the provided statistics, adoption success is best measured by the “number of children who were eligible for adoption on 7/1 who were adopted by 6/30,” which is displayed in Figure 4. In this case, eligibility refers to children for whom a termination of parental rights order has been granted. We estimate that the children listed on the platform comprise between one-quarter and one-third of all eligible children. The remaining children likely already had a path to adoption identified through a foster parent or relative and would be expected to have a faster path to adoption.
In Table 3, we compare the circuit’s average performance against the statewide average for the four years before implementing the platform and the four years after implementation. Because the monthly case creation peaked after July 1, 2018, as the platform’s usage gradually ramped up in 2018, we disregard the annual report for July 1, 2018, to June 30, 2019, as a transition period. Thus, we compare the mean across annual statistics from July 1, 2014, until June 30, 2018, against July 1, 2019, until June 30, 2023. Before implementation, the agency’s performance was only 3% higher than the statewide average, corresponding to an average ranking among all circuits of 9.25 out of 20. Thus, we expect the statewide AFCARS case data used for benchmarking to be accurate.
Considering the averages over the four years since implementation, the circuit has outperformed statewide averages. We note that the statewide average for the percentage of eligible children adopted decreased compared to before 2018, which could reflect higher acuity in children’s needs or increased difficulties in casework and judicial processes from the COVID-19 pandemic. The mean ratio for the circuit metric compared to the statewide metric increased to 13%. While it is difficult to directly link the percentage of eligible children adopted over a one-year time frame to the outcomes explored in Section 8.2, this indication that the circuit has improved its performance in relation to the state as a whole lends credence to the value of our previous analysis using benchmarks from statewide AFCARS data.
Table 3: Percent of eligible children adopted each year for the circuit that implemented a caseworker-driven search platform compared to state averages.
Before After
Comparison Period 7/1/2014-6/30/2018 7/1/2019-6/30/2023
Circuit Annual Mean % of Children Adopted 57% 57%
Statewide Annual Mean % Adopted 55% 51%
Circuit-Statewide Ratio Mean 103% 113%
Mean Rank (of all 20 circuits) 9.25 7.00
9 Discussion and Conclusion
We are the first to apply market design principles and a formal game theoretical model to study the process by which children in the child welfare system are adopted. We introduced a novel search-and-matching model to compare two competing search paradigms. First, we analyzed the Nash equilibria of the games induced by our model. We show that agents follow novel threshold strategies as best responses, subject to a tie-breaking assumption. Also, Nash equilibria form a non-empty complete lattice in either setting.
Search frictions caused by the nature of the adoption process present one of the biggest obstacles to efficient matching efforts. We found that decreasing wasted search efforts leads to generally better outcomes, which are realized in caseworker-driven search. We explored the nuances of this finding to identify circumstances for which this intuition fails to hold.
We have shown that caseworker-driven search can Pareto dominate family-driven search, but no equilibrium of family-driven search can ever Pareto dominate any caseworker-driven search equilibrium. Despite less wasted search efforts, caseworker-driven search equilibria do not always Pareto dominate those of family-driven search. However, we numerically investigated a wide range of underlying model parameters to observe that caseworker-driven search is better for most agents.
Conversely, we also identified environments in which caseworker-driven search would not be desirable. When agents are very patient and preferences are highly correlated, we have numerically shown that a majority of children are worse off in caseworker-driven search. However, such a level of patience is unlikely to be prevalent in practice. Not only is it in children’s best interest to minimize the time in foster care, but we know from domain experts that most families’ patience is also limited: their wish to adopt a child is often very strong after completing the strenuous process of a home study investigation. We have also shown that in such cases — i.e., if families are sufficiently impatient — all children are unambiguously better off in caseworker-driven search compared to family-driven search, regardless of correlation in preferences. Moreover, agents’ preferences are far from being highly correlated in practice. Domain experts report idiosyncratic preferences of families based on different characteristics of children such as ethnicity, gender, or age.
These results suggest that adoption agencies should consider using search technologies that allow caseworkers to perform a targeted search, rather than a strategy of broad announcements to which families respond. This is validated in our case study of an adoption agency in Florida, which, since switching to caseworker-driven search, has outperformed a state-wide benchmark for adoptions at the three-year mark by $17\%$. Our findings show the potential for replacing the main search paradigm in adoption, family-driven search, with newer caseworker-driven search paradigm.
Our work also motivates future empirical research to understand the operations of a child welfare matching system at a granular level. Little is known about the efficacy of different search methods and how caseworkers employ them. Understanding how caseworkers perform their jobs and how technology can help them becomes increasingly important as the workforce faces challenges of burnout and turnover. Furthermore, studying families’ interactions with the platform and the trajectories of their engagement over time represent additional important subjects of analysis. Finally, studying how different search practices affect children who participate in them represents another critical way to increase our understanding of how to serve and protect this vulnerable population.
A Additional Lemmas and Propositions
A.1 Proposition A.1
@@ A Additional Lemmas and Propositions
B Remaining Proofs
B.1 Proof of Proposition 3.1
@@ B Remaining Proofs
$u^{FS*}_{c}(s_{-f})=\delta_{F}u^{FS*}_{c}(s_{-f})+\frac{1}{n}\sum_{c\in M_{f}(s)}\Big{(}\beta_{cf}(s)p\big{(}v_{f}({c})-\delta_{F}u^{FS*}_{c}(s_{-f})\big{)}-\kappa_{F}\Big{)}.$ (17)
@@
and otherwise $s_{c}$ would not be a best response to $s_{-c}$. Again, by our tie-breaking assumption the claim of the proposition follows for children in FS.
B.3 Proof of Proposition 4.5
@@
The inequality holds because of the following reason: Suppose a family $f$ is interested in $c$ under $s(y)$ but not under $s(y^{\prime})$. Since $f$ is weakly less selective in $s(y^{\prime})$, it must be the case that there exists another family $f^{\prime}$ with $v_{c}({f^{\prime}})>v_{c}({f})$ that is not mutually interested in $c$ under $s(y)$ but under $s(y^{\prime})$. Note that for any family that loses interest in $c$ under $s(y^{\prime})$, there must exist such a unique family that replaces it and is preferred by $c$.
@@
B.4 Proof of Lemma 5.2
Proof B.1
@@
B.5 Proof of Theorem 5.5
Proof B.2
@@
Assume that $p\geq\max\{\kappa_{C},\kappa_{F}\}$, and $\epsilon$ and $\delta_{C}$ are sufficiently small. If $p(1-\lambda p)\geq\kappa_{F}$ then the matching correspondences of the unique FSE and the unique CSE are identical, and $f_{2}$ only incurs search costs in CS if the match between $f_{1}$ and $c_{1}$ is not suitable. Thus, an agent can be made strictly better off in CS compared to FS without making anyone else worse off by saving wasted search efforts. If $p(1-\lambda p)<\kappa_{F}$, then in the unique FSE there is only mutual interest between $c$ and $f_{1}$ while in the unique CSE, all agents have again mutually in each other. Furthermore, the CSE is a Pareto improvement over the FSE.
@@
B.6 Proof of Proposition 5.6
Proof B.3
@@
B.7 Proof of Proposition 5.7
Proof B.4
@@
If $\epsilon>0$ is small enough and $p(1-\lambda p)<\kappa_{F}$, then in the unique FSE $s^{FS}$, only $c_{1}$ and $f_{1}$ will be mutually interested in each other and $c_{2}$ and $f_{2}$. Family $f_{2}$ will not be interested in $c_{1}$ in FS independent of how patient $f_{2}$ is. That is, because the probability of actually matching with $c_{1}$ while facing competition from $c_{1}$ is too small, yet $f_{2}$ would have to incur search costs every time $c_{1}$ is active. However, if $\delta_{F}$ is sufficiently large, then the unique CSE matching correspondence is $M(s^{CS})=\{(c_{1},f_{1}),(c_{1},f_{2})\}$. In CS, $c_{2}$, the “low-type” child, will remain unmatched.
B.8 Proof of Proposition 5.8
Proof B.5
@@
B.9 Proof of Proposition 6.1
Proof B.6
@@
B.10 Proof of Proposition 6.2
Proof B.7
@@
Choose $p$ and $\kappa_{F}$, such that $p(1-\lambda^{\prime}p)<\kappa_{F}\leq p(1-\lambda p)$. If $\kappa_{C}$, $\delta_{C}$, and $\delta_{F}$ are sufficiently small, child $c$ will be mutually interested in $f_{1}$ and $f_{2}$ in the unique FSE for $\lambda$. But in the unique FSE for $\lambda^{\prime}$ only $c$ and $f_{1}$ will be mutually interested in each other. However, if the difference between $\lambda$ and $\lambda^{\prime}$ is very small, $c$’s utility can be smaller in the case where the market thickness indicator is $\lambda$, as the increased probability of $f_{1}$ being active might not compensate for the loss of $f_{2}$’s interest.
B.11 Proof of Proposition 6.3
@@
C Algorithm to Compute FS-TSs from Threshold Profiles

### v4

@@ 9.2 Benchmark Against Statewide Outcomes
Figure 5: Actual adoptions by the agency using the platform and other channels compared to Florida AFCARS benchmark model. Dashed lines represent the 95% confidence interval defined by the Poisson binomial distribution.
9.3 Platform Effect in Cox Proportional Hazards Model
To provide additional evidence of the platform’s effectiveness compared to statewide outcomes, we re-estimated the Cox proportional hazards model on a combined dataset to verify that engagement with the online adoption matching platform speeds the journey from TPR to adoption finalization. Because AFCARS restrictions bar direct record linkage and obscure dates of birth (i.e., to the month of birth), a given child can appear in both treatment and control groups — an approach that biases estimates toward understating, rather than exaggerating, any positive or negative platform effect.
Since there is a significant difference between the TPR date for some children and the date on which they were added to the platform, we treat the platform variable as time-varying: the child’s timeline starts at TPR, with the platform variable’s value switching from zero to one at a point three months after the platform case is created. The three-month delay accounts for the minimum period Florida law requires before an adoption can be finalized after placement. Thus, any adoption within the three months after listing cannot be attributed to the platform’s search.
More precisely, when compared with the Cox proportional hazards model introduced in Equation (13), we add a binary covariate to account for whether the child has been on the platform or not. We make the term $\textit{OnPlatform}_{i}(t)$ time dependent, and the binary variable $\textit{OnPlatform}_{i}(t)$ is turned on three months after the child is listed on the platform. Each case’s timeline still begins on the TPR date. [^13] We provide an alternative analysis with non-time-varying platform variables, that is, where the timeline for platform children begins once they are listed on the platform, in online Appendix I.2.
We call this Model 2, with resulting coefficients found in Table 2. We see that control variables behave as expected. Girls experience higher adoption hazards than boys, while Black and Hispanic children tend to be adopted more slowly than children of other racial groups. A clinical disability diagnosis and a child’s increasing age are also linked to decreased adoption hazards, as well as TPR in pandemic years or immediately before.
This analysis finds that the platform is consistently associated with faster adoption. The estimated platform hazard ratio is 1.272, indicating a 27.2% increase in the hazard rates compared to not using the platform (i.e., faster adoptions on average). This ratio is significant at the 1% level, despite the biases against the platform built into the dataset construction mentioned above. Taken together, these results confirm that the agency has seen accelerated adoption and finalization rates compared to statewide outcomes after implementing the platform.
Conclusion
Treating the search for adoptive families as an operations challenge, we develop the first formal game-theoretic model for the child welfare system adoption process and introduce a novel search-and-matching framework to compare caseworker-driven and family-driven search paradigms. We characterize the Nash equilibria of these models, showing that agents adopt threshold strategies under a mild tie-breaking condition and that equilibria form a non-empty complete lattice. We find that caseworker-driven search better avoids wasted search efforts and, in most settings, improves outcomes. While caseworker-driven equilibria can Pareto dominate all family-driven equilibria, the converse does not hold. Although caseworker-driven equilibria do not always Pareto dominate those of family-driven search, they typically perform better for most agents, as confirmed through extensive numerical analysis. However, we also identify conditions — notably, highly patient agents with strongly correlated preferences — under which caseworker-driven search may harm children. Yet such conditions are rare in practice: families usually desire a timely placement after completing the adoption approval process, and preferences tend to be idiosyncratic. When families are sufficiently impatient, caseworker-driven search unambiguously benefits all children.
Empirically, we analyze adoption outcomes for 279 children in need of adoptive placements who were served by a Florida agency that shifted from family-driven to caseworker-driven search. Using Cox proportional hazards models that control for observable characteristics as a benchmark, we find a 24% increase in the likelihood of adoption finalization within three years. Similarly, a hazard model fitted on a combined dataset indicates 27% increases in the instantaneous adoption rate despite conservative bias in the data construction.
Our work provides a foundation for future empirical research into managing the operation of adoptive placement searches, especially how to allocate caseworkers’ efforts across difference search channels. While existing research in child welfare literature emphasizes the effectiveness of highly specialized recruiters, we show how technology can help frontline caseworkers achieve better outcomes for children. Furthermore, our benchmarking approach that utilizes AFCARS data provides a replicable tool for evaluating agency performance and child welfare system reforms. Additional study of caseworker behavior, family engagement, and children’s outcomes under different search practices is essential to improving services for this vulnerable population.
A Home Studies
Before a family can be considered for adoption, they have to complete a home study. State regulations determine minimum requirements for home studies. For example, 2024 Florida Statutes Chapter 63 Section 092 states:
The preliminary home study must be made to determine the suitability of the intended adoptive parents and may be completed prior to the identification of a prospective adoptive child. The study must include, at a minimum, the following:
  • An interview with the intended adoptive parents.
  • Records checks of the department’s central abuse registry, which the department shall provide to the entity conducting the preliminary home study, and criminal records correspondence checks under s. 39.0138 through the Department of Law Enforcement on the intended adoptive parents.
  • An assessment of the physical environment of the home.
  • A determination of the financial security of the intended adoptive parents.
  • Documentation of counseling and education of the intended adoptive parents on adoptive parenting, as determined by the entity conducting the preliminary home study. The training specified in s. 409.175(14) shall only be required for persons who adopt children from the department.
  • Documentation that information on adoption and the adoption process has been provided to the intended adoptive parents.
  • Documentation that information on support services available in the community has been provided to the intended adoptive parents.
  • A copy of each signed acknowledgment of receipt of disclosure required by s. 63.085.
However, most agencies perform more comprehensive studies than required by law to ensure a smoother process. In practice, home studies typically also include additional information on
  • Family background, including education and health
  • Relationships and social environment
  • Parenting experiences
  • A description of daily life routines
  • Details about home and neighborhood
  • Reasons for seeking an adoptive placement
  • Recommendations by the caseworker about what types of children the family is suited for (e.g., especially regarding special needs, but also other factors such as gender or age)
B Increased Uncertainty or Heterogeneity
Our analytical model assumes a relatively limited amount of variability and uncertainty: the value of every match is known a priori. The only uncertainty is whether the child is compatible with the considered family, with the suitability probability $p$ being the same for all matches. This is not without loss of generality. In practice, compatibility is a spectrum, and values may be uncertain. This means that while caseworkers and families may have prior beliefs based on their own experiences and potentially supplemented by recommender systems, the actual value of a match is not known until the investigation is conducted. Unfortunately, explicitly modeling such higher degrees of uncertainty is intractable in a model where both market sides are strategic and more than one match may be investigated per time period.
This is because if values are uncertain, family utility is no longer monotonic in terms of how selective a child is. While children being less selective (i.e., considering families it has lower expected value for) still increases a family’s utility if it causes the child to be interested in them, it often reduces a family’s utility if the child was already interested in them (as the child has some probability of instead matching with a different family they didn’t even consider before). This may lead to non-convergence of best responses and non-existence of equilibria in pure strategy, even if only either the suitability or market thickness is heterogeneous.
Proposition B.1
If suitability probabilities are heterogeneous, i.e., if there exist different $p_{c,f}$ for different child/family pairs, or if families have heterogeneous probabilities $\lambda_{f}$ to be present, then pure strategy equilibria do not always exist.
Proof B.2
Proof. Consider the following example in FS where each child/family pair has its own probability $p_{c,f}$ for being suitable. Let $C=\{c_{1},c_{2}\}$, $F=\{f_{1},f_{2},f_{3}\}$, and let valuations and suitability probabilities be according to the following tables
$v_{c}({f})$ $f_{1}$ $f_{2}$ $f_{3}$
$c_{1}$ $2$ $1$ $0.5$
$c_{2}$ $0.5$ 00 $1$
$v_{f}({c})$ $f_{1}$ $f_{2}$ $f_{3}$
$c_{1}$ $20$ $0.11$ $2$
$c_{2}$ $5$ 00 $2$
$p_{c,f}$ $f_{1}$ $f_{2}$ $f_{3}$
$c_{1}$ $0.1$ $0.99$ $1$
$c_{2}$ $0.5$ $0.5$ $0.5$
Let $\lambda=1$ (i.e., all family types are always present). Further, assume children are very impatient and have negligible search costs, i.e. $\delta_{C}=\kappa_{C}=0$, while families are very patient $\delta_{F}=1$ and have search cost $\kappa_{F}=0.1$. Note that this implies that a) both children are so impatient, that they will be interested in all families and b) families will be so patient, that they are at most interested in a single-child type (unless both children types give almost the same expected utility, which cannot happen in pure strategy with these parameters). Note that setting these extreme values is not required for equilibrium non-existence and is only done for brevity.
In this setting, it is not rational for $f_{2}$ to be interested in $c_{2}$, independent of the other agent’s strategies. Further, if $f_{1}$ is interested in $c_{1}$ in equilibrium, $f_{2}$ cannot be interested in $c_{1}$ as doing so would result in negative utility, as $c_{1}$ prefers $f_{1}$ to $f_{2}$: the probability that $c_{1}$ doesn’t successfully match with $f_{1}$ times the probability that $f_{2}$ and $c_{1}$ are compatible times $v_{f_{2}}(c_{1})$ equals $(1-0.1)\times 0.99\times 0.11=0.09801\leq 0.1=\kappa_{F}$. However, if $f_{1}$ is not interested in $c_{1}$, $f_{2}$ would obtain positive utility (i.e., $0.99\times 0.11=0.1089\geq 0.1$) and therefore would be interested.
Thus, we only have 4 possible equilibrium candidates, depending on in which child families $f_{1}$ and $f_{3}$ are interested.
  1. 1.
    If $f_{1}$ and $f_{3}$ are interested in $c_{1}$ in equilibrium, then $f_{1}$ will obtain utility $0.1\times 20-0.1=1.9$. However, $f_{1}$ deviating to be (solely) interested in $c_{2}$ would yield utility $0.5\times 5-0.1=2.4>1.9$, a contradiction.
  2. 2.
    If $f_{1}$ is interested in $c_{1}$ and $f_{3}$ is interested in $c_{2}$ in equilibrium, then $f_{3}$ will obtain utility $0.5\times 2-0.1=0.9$. However, $f_{3}$ deviating to be (solely) interested in $c_{1}$ would yield utility $(1-0.1)\times 1\times 2-0.1=1.7>0.9$, a contradiction.
  3. 3.
    If $f_{1}$ and $f_{3}$ are interested in $c_{2}$ in equilibrium, then $f_{1}$ will only be chosen if $f_{3}$ is incompatible, thus obtaining expected utility $0.5\times 0.5\times 5-0.1=1.15$. However, $f_{1}$ deviating to be (solely) interested in $c_{1}$ would yield utility $0.1\times 20-0.1=1.9>1.15$, a contradiction.
  4. 4.
    If $f_{1}$ is interested in $c_{2}$ and $f_{3}$ is interested in $c_{1}$ in equilibrium, then $f_{2}$ is interested in $c_{1}$. Since $c_{1}$ prefers $f_{2}$ and has a $p_{c_{1},f_{2}}=0.99$ likelihood of being compatible, this implies that $f_{3}$ only obtains utility $(1-0.99)\times 1\times 2-0.1=-0.08$, an immediate contradiction.
In conclusion, there can be no equilibrium in pure strategies, because $f_{1}$ always prefers the child type $f_{3}$ is not interested in, while $f_{3}$ always prefers the child type $f_{1}$ is interested in.
To see the nonexistence of equilibria for heterogeneous $\lambda_{f}$, consider the following example. Let $C=\{c_{1},c_{2}\}$, $F=\{f_{1},f_{2},f_{3}\}$ and let valuations and $\lambda_{f}$ be according to the following tables
$v_{c}({f})$ $f_{1}$ $f_{2}$ $f_{3}$
$c_{1}$ $2$ $1$ $0.5$
$c_{2}$ $0.5$ 00 $1$
$v_{f}({c})$ $f_{1}$ $f_{2}$ $f_{3}$
$c_{1}$ $4$ $0.11$ $2$
$c_{2}$ $5$ 00 $1$
$f_{1}$ $f_{2}$ $f_{3}$
$\lambda_{f}$ $0.1$ $0.99$ $1$
Let $p=1$ (i.e., all families are suitable). Further, assume children are impatient and have negligible search costs, i.e. $\delta_{C}=\kappa_{C}=0$, while families are very patient $\delta_{F}=1$ and have search cost $\kappa_{F}=0.1$. Note that this implies that a) both children are so impatient, that they will be interested in all families and b) families will be so patient, that they are at most interested in a single-child type (unless both children types give almost the same expected utility, which cannot happen in pure strategy with these parameters). Note that setting these extreme values are not required for equilibrium non-existence and is only done for brevity.
In this setting, it is not rational for $f_{2}$ to be interested in $c_{2}$, independent of the other agent’s strategies. Further, if $f_{1}$ is mutually interested in $c_{1}$ in equilibrium, $f_{2}$ cannot be interested in $c_{1}$ as doing so would result in negative utility: due to $c_{1}$ preferring $f_{1}$ whenever present ($10\%$ of the time), the immediate expected value of showing interest is only $(1-0.1)\times 0.11=0.099$, less than the search cost of $0.1$. However, if $f_{1}$ is not mutually interested in $c_{1}$, $f_{2}$ is the child’s first choice and would obtain positive utility (i.e., $0.11\geq 0.1$). Therefore, $f_{2}$ would be interested.
Thus, we only have 4 possible equilibrium candidates, depending on which child families $f_{1}$ and $f_{3}$ are interested in.
  1. 1.
    If $f_{1}$ and $f_{3}$ are interested in $c_{1}$ in equilibrium, $f_{1}$ will obtain utility $4-0.1=3.9$. However, $f_{1}$ deviating to be (solely) interested in $c_{2}$ would yield utility $5-0.1=4.9>3.9$, a contradiction.
  2. 2.
    If $f_{1}$ is interested in $c_{1}$ and $f_{3}$ is interested in $c_{2}$ in equilibrium, $f_{3}$ will obtain utility $1-0.1=0.9$. However, $f_{3}$ deviating to be (solely) interested in $c_{1}$, it would successfully match whenever $f_{1}$ is not present, yielding utility $(1-0.1)\times 2-0.1=1.7>0.9$, a contradiction.
  3. 3.
    If $f_{1}$ and $f_{3}$ are interested in $c_{2}$ in equilibrium, $f_{1}$ will obtain no utility, as $c_{2}$ will always match successfully with $f_{3}$, a contradiction since $f_{1}$ can guarantee positive utility by being interested in $c_{1}$.
  4. 4.
    If $f_{1}$ is interested in $c_{2}$ and $f_{3}$ is mutually interested in $c_{1}$ in equilibrium, $f_{2}$ is interested in $c_{1}$. Since $c_{1}$ prefers $f_{2}$ over $f_{3}$, $f_{3}$ only obtains utility $(1-0.99)\times 1\times 2-0.1=-0.08$, an immediate contradiction.
In conclusion, there can be no equilibrium in pure strategies, because $f_{1}$ always prefers the child type $f_{3}$ is not interested in, while $f_{3}$ always prefers the child type $f_{1}$ is interested in.
[…]
As the same effect can be recreated by uncertainty about match values, we immediately obtain the following.
Corollary B.3
If values are uncertain, pure strategy equilibria do not always exist.
While equilibria still exist in mixed strategy (i.e., where agents play randomized strategies), it is intractable to fully characterize or calculate mixed strategy equilibria in such a complex system.
While we, therefore, restrict our formal analysis to a low level of uncertainty, it is important to note that uncertainty typically favors CS: In CS, the caseworker can decide whether to investigate another family based on the expected utility gain while already knowing the true value of previously investigated matches. Additionally, investigating a family does not prevent the caseworker from going back and matching with previously investigated families. This allows caseworkers more flexibility than FS, where all investigation decisions must be made simultaneously, increasing the chance of “wasted” investigations.
However, our insight that neither approach always dominates the other still holds. Similar to before, the lower cost of being mutually interested in CS changes strategic considerations on both sides of the market and can lead to more matches in FS (e.g., if families are so patient that they are only interested in a small set of very attractive children in CS).
C Additional Lemmas and Propositions
C.1 Proposition C.1
@@ C Additional Lemmas and Propositions
D Remaining Proofs
D.1 Proof of Proposition 4.1
@@ D Remaining Proofs
$u^{FS*}_{f}(s_{-f})=\delta_{F}u^{FS*}_{f}(s_{-f})+\frac{1}{n}\sum_{c\in M_{f}(s)}\Big{(}\beta_{cf}(s)p\big{(}v_{f}({c})-\delta_{F}u^{FS*}_{f}(s_{-f})\big{)}-\kappa_{F}\Big{)}.$ (19)
@@
and otherwise $s_{c}$ would not be a best response to $s_{-c}$. Again, by our tie-breaking assumption, the claim of the proposition follows for children in FS.
D.3 Proof of Proposition 5.5
@@
The inequality holds for the following reason: Suppose a family $f$ is interested in $c$ under $s(y)$ but not under $s(y^{\prime})$. Since $f$ is weakly less selective in $s(y^{\prime})$, it must be the case that there exists another family $f^{\prime}$ with $v_{c}({f^{\prime}})>v_{c}({f})$ that is not mutually interested in $c$ under $s(y)$ but under $s(y^{\prime})$. Note that for any family that loses interest in $c$ under $s(y^{\prime})$, there must exist such a unique family that replaces it and is preferred by $c$.
@@
D.4 Proof of Lemma 6.2
Proof D.1
@@
D.5 Proof of Lemma 6.3.
Proof D.2
D.6 Proof of Theorem 6.4
Proof D.3
@@
Assume that $p\geq\max\{\kappa_{C},\kappa_{F}\}$, and $\epsilon$ and $\delta_{C}$ are sufficiently small. If $p(1-\lambda p)\geq\kappa_{F}$ then the matching correspondences of the unique FSE and the unique CSE are identical, and $f_{2}$ only incurs search costs in CS if the match between $f_{1}$ and $c_{1}$ is not suitable. Thus, an agent can be made strictly better off in CS compared to FS without making anyone else worse off by saving wasted search efforts. If $p(1-\lambda p)<\kappa_{F}$, then in the unique FSE there is only mutual interest between $c$ and $f_{1}$ while in the unique CSE, all agents have again mutual interest in each other. Furthermore, the CSE is a Pareto improvement over the FSE.
@@
D.7 Proof of Proposition 6.5
Proof D.4
@@
D.8 Proof of Proposition 6.6
Proof D.5
@@
If $\epsilon>0$ is small enough and $p(1-\lambda p)<\kappa_{F}$, then in the unique FSE $s^{FS}$, only $c_{1}$ and $f_{1}$ will be mutually interested in each other and $c_{2}$ and $f_{2}$. Family $f_{2}$ will not be interested in $c_{1}$ in FS, independent of how patient $f_{2}$ is. That is, because the probability of actually matching with $c_{1}$ while facing competition from $c_{1}$ is too small, yet $f_{2}$ would have to incur search costs every time $c_{1}$ is active. However, if $\delta_{F}$ is sufficiently large, then the unique CSE matching correspondence is $M(s^{CS})=\{(c_{1},f_{1}),(c_{1},f_{2})\}$. In CS, $c_{2}$, the “low-type” child, will remain unmatched.
D.9 Proof of Proposition 6.7
Proof D.6
@@
D.10 Proof of Proposition 7.1
Proof D.7
@@
D.11 Proof of Proposition 7.2
Proof D.8
@@
Choose $p$ and $\kappa_{F}$, such that $p(1-\lambda^{\prime}p)<\kappa_{F}\leq p(1-\lambda p)$. If $\kappa_{C}$, $\delta_{C}$, and $\delta_{F}$ are sufficiently small, child $c$ will be mutually interested in $f_{1}$ and $f_{2}$ in the unique FSE for $\lambda$. But in the unique FSE for $\lambda^{\prime}$, only $c$ and $f_{1}$ will be mutually interested in each other. However, if the difference between $\lambda$ and $\lambda^{\prime}$ is very small, $c$’s utility can be smaller in the case where the market thickness indicator is $\lambda$, as the increased probability of $f_{1}$ being active might not compensate for the loss of $f_{2}$’s interest.
D.12 Proof of Proposition 7.3
@@
E Algorithm to Compute FS-TSs from Threshold Profiles